Assume that a sample is used to estimate a population proportion p. Find the 90% confidence interval for a sample of size 279 with 38% successes. Enter your answer as an open - interval (i.e., parenthesis) using decimals (not percents) accurate to three decimal places.

Confidence Interval = ______________.
Answer #1

n = 279

\hat{p} = 0.38

\alpha = 0.10

From Table, criticalvalues of Z = \pm 1.645

Confidence Interval:

0.38\pm 1.645\times \sqrt{\frac{0.38\times (1-0.38)}{279}}=(0.332,0.428)

So,

Answer is:

(0.332, 0.428)